Please help: The sum of two numbers is 17. If the larger is two more than twice the smaller, the larger number.
Answer by nerdybill(7384)

Please help: The sum of two numbers is 17. If the larger is two more than twice the smaller, the larger number.
.
Let x = smaller integer
then because "the larger is two more than twice the smaller"
2x+2 = larger integer
.
x + (2x+2) = 17
3x+2 = 17
3x = 15
x = 5 (smaller integer)
.
larger integer:
2x+2 = 2(5)+2 = 10+2 = 12 (larger integer)